Determine the focal length of a thin converging lens if the power is 5.0 dioptres
1
Expert's answer
2019-09-10T13:41:24-0400

Optical power of a lens is given by formula

"P = \frac{1}{f}," where f is the focal length.

Reciprocally,

"f = \frac{1}{P} = \frac{1}{5.0\text{ dpt}} = 0.2\text{ m} = 20\text{ cm}."